OpenStudy (anonymous):

When Y = 50, what is X in the function, √100 ·X - 50 = Y?

OpenStudy (anonymous):

i think its 90

OpenStudy (ajanijones):

X = 1

OpenStudy (ajanijones):

Medal please.

OpenStudy (anonymous):

\[\sqrt{100} * X -50 =50\]

OpenStudy (anonymous):

+50 on both sides

OpenStudy (ajanijones):

Its not Division its 100 * 1-50 =50

OpenStudy (anonymous):

i know but theres a square root in there

OpenStudy (anonymous):

square root of 100 is 10 and subtract that on both sides to get x=90

OpenStudy (anonymous):

actually its divide by 10 on both sides not subtract

OpenStudy (anonymous):

dividing by 10 will give u x=10

OpenStudy (ajanijones):

No you dont divide.